Please help me I tried this but the answers I came up with are
x=-19
y=-17
z=44
Solve the system of equations:
2x−5y+z=−11
−x+y−z=1
x−3y=−7



Answer by Boreal(15235)

2x−5y+z=−11
−x+y−z=1
----------------set x in terms of y and substitute into the other two equations. Now there are two variables in two equations.
x−3y=−7
x=3y-7
2(3y-7)-5y+z= -11, or 6y-14-5y+z=-11 or y+z=3.
-3y+7+y-z=-1 or -2y-z=-8
-------------------
y+z=3
-2y-z=-8
-y=-5
y=5
z= -2
Put into third to get x. -x+5+2=1, and -x=-6, x=6
check in first
2(6)-5(5)+2=12-25+2=-11
{6,5,-2}